Use the following data and the high-low method to determine total fixed costs:

Units Total
Produced Costs
January 1000 $45,550
February 1500 $52,000
March 2100 $61,500
April 1800 $62,000
May 750 $41,250
Answer #1

Variable cost per unit

= (Cost at highest activity - Cost at lowest activity)/(Highest activity-Lowest activity)

= (61,500-41,250)/(2100-750) = 20,250/1350

= $15 per unit

Total Cost = Total Variable cost + Fixed cost

61,500 = (15*2100) + Fixed cost

Fixed cost = 61,500 - 31,500

= $30,000